OpenStudy (amenah8):

WILL MEDAL Optimizaiton: An apartment complex on Ferenginar with 250 units currently has 170 occupants. The current rent for a unit is 510 slips of Gold-Pressed Latinum. The owner of the complex knows from experience that he loses one occupant every time he raises the rent by 1.5 slips of Latinum. Since "profit is its own reward", the owner wants to maximize his profit so he asks for our help, even though he knows that "free advice is seldom cheap".

OpenStudy (amenah8):

the derivative is 2*(2)x1, or 4x, but what do i do from here?

OpenStudy (518nad):

occupants = (510-rent)/1.5 + 170 profit = rent*occupants maximize profit

OpenStudy (amenah8):

now i have to take the derivative of that function, right?

OpenStudy (518nad):

do u understand how occupants and profit equations came about

OpenStudy (518nad):

there are many ways to rewrite this, but the final profit equation will be equivalent

OpenStudy (amenah8):

i think i understand how you got the function

OpenStudy (518nad):

okay so from the 2 equations, can you write, profit as a function of just rent or just occupants

OpenStudy (amenah8):

P(O) = (510-(rent*occupant))/(1.5+170)???? (510-rent)/1.5 + 170 profit = rent*occupants

OpenStudy (amenah8):

i'm not sure if that's right